2x+2y-z=10
X+y+z=11
2x+y+-z=7
So x,y,and z =?

Answer by Alan3354(69443)

2x+2y-z=10
X+y+z=11
2x+y+-z=7
-----------
Since all z terms have a coefficient of 1, I would use elimination.
---
Add Eqn 1 and 2:
2x+2y-z=10
X+y+z=11
------------------- Add
2x + 3y = 21 Eqn A
---
Add Eqn 2 & 3
X+y+z=11
2x+y+-z=7
-------------- Add
3x + 2y = 18 Eqn B
---
4x + 6y = 42 Eqn A times 2
9x + 6y = 54 Eqn B times 3
---------------------------------- Subtract
-5x = -12
x = 12/5
----------
Sub for x in Eqn A or B to find y
Sub for x & y in any Eqn to find z